Kl(aq) + H2O(l) + KMnO4(aq) 166.0/mol 18.02g/mol 153.03g/mol l2(s) + MnO2(s) + KOH(aq) 253.80g/mol 86.97g/mol 56.11g/mol A.) Calculate the mass of manganese (lV) oxide that can be synthesized from 15.00 grams of potassium iodide. B.) Calculate the percent yield for this experiment if a mass of 1.982 grams of manganese (lV) oxide is produced. C.) Determine the percent error D.) compute for the excess amount of the excess reactant
1
Expert's answer
2021-10-27T06:41:36-0400

6KI(aq) +4H2O(l)+2KMnO4(aq)"\to" 3l2(s) +2 MnO2(s)+8KOH(aq)


REM of KI=166.00 g/mol

Moles of KI=15/ 166

= 0.09036moles

Mole ratio= KI:MnO2 = 6:2 = 3:1

Moles of MnO2= 0.09036/3

=0.03012moles

REM of MnO2= 86.99 g/mol

Mass of MnO2= 0.03012× 86.97

=2.6195364g




Percent Yield =44.47%